how do I find the standard curve that lies to the right.
(a)z=1.00 (b)z=-1.05 (c) z=0 (d)z=2.80 (e)z=1.96

Answer by ikleyn(52879)

.
how do I find the standard curve that lies to the right.
(a)z=1.00 (b)z=-1.05 (c) z=0 (d)z=2.80 (e)z=1.96
----------------------------------------------------------

What do you mean ?

Read your post.